# Efficient Computation of Confidence Sets Using Classification on Equidistributed Grids

### v1

@@ 2.2 Literature Review
Aradillas-Lopez and Gandhi, (2016) develop a structural model of chain-story entry with discrete intensity and obtain the asymptotic properties of the pivotal test statistic via kernel-based estimation, which is then used to construct the confidence set. A grid search is then applied in the parameter space to determine numerically the region of acceptance via the criterion defined by the authors. Ciliberto and Tamer, (2009) study the effect of general firm heterogeneity on the market structure in the context of U.S. airline industry, and propose a moment inequalities-based procedure to conduct inference on the parameters of interest. They employ simulated annealing algorithm to approximate the global minimum of the objective criterion along with subsampling to obtain the critical value (as described in Chernozhukov et al., (2007)) in the construction of the confidence region.
@@ 3.2 Asymptotic Behavior
where ϕ()italic-$\phi(\cdot)$ is implicitly defined through the RBF kernel, i.e. $\phi(\mathbf{x})\cdot\phi(\mathbf{x}^{\prime})=\exp\left(-\frac{||\mathbf{x}-\mathbf{x}^{\prime}||^{2}}{2\sigma^{2}}\right)$, classifies all training examples correctly and classifies the rest of the space as class 0.
@@ 5 Conclusion
In this paper, we devise an efficient and reliable way of reproducing the confidence sets in the context of moment inequalities. These problems typically involve inverting the test and sub-sampling or re-sampling when constructing critical values, and are computationally complex especially when the dimensionalily of the parameter grows. Combining the uses of the support vector machine and the equidistributed sequences, we are able to consistently re-produce the limiting confidence regions as the sample size grows. This method is very easy to implement and provides much better efficiency than the exhaustive search, under very common assumptions and mild conditions. Researchers can generate grid points and fit the SVM classifier utilizing the existing packages in standard programming languages. Researchers can also incorporate an iterative procedure into the generation of grid points for better approximation of the finite-sample confidence sets. This approach might also be used to compute and characterize the identified sets, or the confidence sets of the identified sets, so long as they have a similar formulation.
@@ Appendix A Equidistributed Sequences
Now we present how the grid points look under different generating sequences. Figures 7 through 10 show in the two-dimensional space the grid points generated from Monte Carlo sequences, Sobol sequences, Weyl sequences, Bakers sequences, respectively. Going from the top-left panel to the bottom-right, we present the grids of 20, 100, 500, and 2000 grid points in each of the following four figures.
@@ B.2 Proof of Theorem 1
By Lemma 2, $\forall\theta\in\text{int}(CS)$, $\exists N\in\mathbb{N}$ such that $\forall n>N$, $||I_{n}(\theta)-\theta||^{2}<||E_{n}(\theta)-\theta||^{2}$, and hence such $\sigma_{n}^{2}$ exists that satisfies 21. Therefore, choosing the tuning parameter this way for each $n>N$ makes the influence of the nearest interior grid dominate the influence of all exterior grid points in the limit. Consequently, the SVM classifier trained over these grids under such tuning classifies any $\theta\in CS$ to have label 1; that is, $\{\theta\in CS\}\Rightarrow\{\theta\in SVM_{+}\}$. Because $\mathbb{P}(\theta_{0}\in CS)\geq 1-\alpha$, it follows that
@@ B.3 Proof of Corollary 1
Since for large enough sample size $n$, $||I_{n}(\tilde{\theta})-\tilde{\theta}||^{2}>||E_{n}(\tilde{\theta})-\tilde{\theta}||^{2}$, the chain of inequalities in 22 is well-defined. Therefore, we can find values for the tuning parameter $\sigma_{n}^{2}$ for all $n$’s that are sufficiently large such that a point $\tilde{\theta}$ in the complement of $CS$ will be labelled $-1$ by the SVM classifier.
Combining this with Theorem 1, it follows that when the sample size is large enough, by taking the values in the intersection of 21 and 22 for the tuning parameter $\sigma^{2}$, we have that a point $\theta\in\Theta$ belongs to $CS$ if and only if the SVM classifier under such tuning classifies it to have label $1$. ∎

### v2

@@ 2.2 Literature Review
Aradillas-Lopez and Gandhi, (2016) develop a structural model of chain-story entry with discrete intensity and obtain the asymptotic properties of the pivotal test statistic via kernel-based estimation, which is then used to construct the confidence set. A grid search is then applied in the parameter space to determine numerically the region of acceptance via the criterion defined by the authors. Ciliberto and Tamer, (2009) study the effect of general firm heterogeneity on the market structure in the context of U.S. airline industry, and propose a moment inequalities-based procedure to conduct inference on the parameters of interest. They employ simulated annealing algorithm to approximate the global minimum of the objective criterion along with subsampling to obtain the critical value (as described in Chernozhukov et al., (2007)) in the construction of the confidence region. We discuss more details of their approach in Subsection 4.3.
@@ 3.2 Asymptotic Behavior
where $\phi(\cdot)$ is implicitly defined through the RBF kernel, i.e. $\phi(\mathbf{x})\cdot\phi(\mathbf{x}^{\prime})=\exp\left(-\frac{||\mathbf{x}-\mathbf{x}^{\prime}||^{2}}{2\sigma^{2}}\right)$, classifies all training examples correctly and classifies the rest of the space as class 0.
@@ 4.2 OLS Coefficients
4.3 Moment Condition Models
We simulate the construction of the confidence set based on moment conditions in this subsection. In particular, we follow the approach proposed in Chernozhukov et al., (2007) (henceforth CHT) and employed in Ciliberto and Tamer, (2009) where the confidence set is based on keeping the all parameter vectors that satisfy an objective function resulted from moment conditions. That is, the collection of parameter values the evaluation of the criterion at which meets the moment conditions, i.e.
$C_{n}(c)=\{\theta\in\Theta:n[\hat{Q}_{n}(\theta)-\min_{t}\hat{Q}(t)]\leq c\}$ (21)
Note that this approach involves minimization of the objective function for the entire data followed by minimization for each bootstrap subsample using some off-the-shelf solver, such as Nealder-Mead and other genetic algorithm like simulated annealing in Ciliberto and Tamer, (2009). Subsequently, a grid search over $C_{n}(c_{(0)})$ given some initial $c_{(0)}$ is performed to find the maximum of $b(Q_{b}(\theta)-\min_{t}Q_{b}(t))$ for each bootstrap sample of size $b$. Once all bootstrap optimal values are computed, the empirical distribution of the values is formed and the empirical $\alpha$-quantile is chosen to replace $c$ in 7. Ciliberto and Tamer, (2009) repeats the above construction of critical value $c$ one more time, and then retains all qualified parameter vectors via another grid search.
Consider a standard least squares problem with 3 parameter of interest, an intercept and two slope coefficients, as the objective. We set the sample size $n=1,000$ and bootstrap sample size $b=250$, and draw some $X\in\mathbb{R}^{3}$ to make $Y$ for each observation. For simplicity, we assume the model is correctly specified and hence obviating the inner minimization in the the above construct (7), which reduces to the set $\{\theta\in\Theta:n\hat{Q}_{n}(\theta)\leq c\}$ where $\hat{Q}_{n}(\theta)=\frac{1}{n}\sum_{i}(Y_{i}-X_{i}^{\top}\theta)^{2}$. We use Nealder-Mead solver to obtain the minimal values of $\hat{Q}_{n}(\theta)$ using the entire sample as well as in the bootstrap sub-routine. The parameter space, a subset of $\mathbb{R}^{3}$, is divided into a uniform grid on the cube consisting of 101 values for each parameter and hence a total of $101^{3}$ grid points. This is arguably a very coarse grid since the points along each parameter dimension are fairly sparsely located, containing only 5 values between each pair of integers. Nonetheless, performing the brute-force CHT routine on such grid, with a programming choice of R without parallelization, runs for 3 hours on Penn State Roar Collab server with 16GB memory. We repeat the same exercise on a smaller grid with 51 points along each dimension and use the larger grid as benchmark to examine the difference between standard procedure and the use of SVM classifier.
Figure 7: The accepted grid points (blue) form a non-rectangular area. Reporting interval for each parameter leads readers to potentially falsely accept points outside of the joint confidence set (red).
The relationship between the two slope coefficients is depicted in Figure 7 based on the search on the $51^{3}$ grid. The blue dots are accepted according to CHT procedure and the red crosses are rejected. We note that once we report the confidence interval for each parameter, information is lost in that both blue dots and red crosses appear to be in the joint rectangle. For the SVM classifier, we train the model based on the coarser grid and examine its ability to capture the underlying confidence region using the large grid of $101^{3}$ points. As described above, labels for the purpose of classification are assigned according to each grid point’s membership to the confidence set, i.e. 1 if the grid point is accepted according to CHT and -1 otherwise. We calculate the percentages of grid points classified to be in the CHT CS out of all grid points on both the training grid (i.e. the $51^{3}$ grid) and the test grid (i.e. the $101^{3}$ grid), denoted as “Labeled CS”. We also compare the predicted labels against the labels dictated by the CHT procedure, and denote the percentage of labels agreed by both approaches as “Correct”. Results are reported in Table 2. The same calculations for the CHT procedure with rectangular region is also included for exposition. We emphasize again that, rather than making wrong predictions, the collection of individual intervals does not tell us anything about the confidence set but its minimum bounding box, which the SVM classifier is able to overcome.
An immediate observation is the rectangle falsely includes many grid points that do not belong to the CS and leads to lower statistical power due to the fact that the joint confidence region does not admit a rectangular shape. To correct this means we would have to store all the accepted grid points; and even so, there is generally no prescription on new parameter vectors, e.g. a point that lies in the convex hull but outside of the concave hull of the accepted grid points, without evaluating the criterion that is based on moment conditions. Therefore, it would be much more concise and precise to consider instead the decision function of the form in equation (20) for $K(\theta_{1},\theta_{2})=\exp\{||\theta_{1}-\theta_{2}||^{2}/\sigma^{2}\}$ with a properly tuned $\sigma^{2}$. This allows us to easily determine if any parameter vector including those outside the grid points belongs to the joint confidence region by simply evaluating the function at the parameter vector. Lastly, we note that while such grid of $51^{3}$ points is fairly coarse, it is already so abundant that tuning of the kernel is essentially unnecessary if time is of utmost important. In the un-tuned classifier, the number of support vectors is 4,901, which amount to summing up 4,901 evaluations of simple operations. It takes about 16 seconds in the aforementioned computing environment for the trained classifier to classify each 10,000 grid points. The properly tuned classifier is comprised of 56,004 support vectors in this case.
Method % Labeled CS ($51^{3}$) % Correct ($51^{3}$) % Labeled CS ($101^{3}$) % Correct ($101^{3}$)
Rectangle 86.275 58.758 87.128 44.695
SVM (no tuning) 45.053 99.928 45.165 98.540
SVM (tuned) 45.032 100.000 44.681 99.024
Table 2: Comparison between standard reporting of intervals and SVM classifiers.
@@ 5 Conclusion
In this paper, we devise an efficient and reliable way of reproducing the confidence sets in the context of moment conditions. These problems typically involve inverting the test and sub-sampling or re-sampling when constructing critical values, and are computationally complex especially when the dimensionalily of the parameter grows. Combining the uses of the support vector machine and the equidistributed sequences, we are able to consistently re-produce the limiting confidence regions as the sample size grows under common assumptions and mild conditions. This procedure is very easy to implement in practice. Researchers can generate grid points and fit the SVM classifier utilizing the existing packages in standard programming languages. Researchers can also incorporate an adaptive procedure into the generation of grid points for better approximation of the confidence sets when the sample size is fixed. Additionally, the use of decision functions provides much higher efficiency and precision than reporting the rectangular regions from the exhaustive search. This approach might also be used to compute and characterize the identified sets, or the confidence sets of the identified sets, so long as they have a similar formulation.
@@ Appendix A Equidistributed Sequences
Now we present how the grid points look under different generating sequences. Figures 8 through 11 show in the two-dimensional space the grid points generated from Monte Carlo sequences, Sobol sequences, Weyl sequences, Bakers sequences, respectively. Going from the top-left panel to the bottom-right, we present the grids of 20, 100, 500, and 2000 grid points in each of the following four figures.
@@ B.2 Proof of Theorem 1
By Lemma 2, $\forall\theta\in\text{int}(CS)$, $\exists N\in\mathbb{N}$ such that $\forall n>N$, $||I_{n}(\theta)-\theta||^{2}<||E_{n}(\theta)-\theta||^{2}$, and hence such $\sigma_{n}^{2}$ exists that satisfies 22. Therefore, choosing the tuning parameter this way for each $n>N$ makes the influence of the nearest interior grid dominate the influence of all exterior grid points in the limit. Consequently, the SVM classifier trained over these grids under such tuning classifies any $\theta\in CS$ to have label 1; that is, $\{\theta\in CS\}\Rightarrow\{\theta\in SVM_{+}\}$. Because $\mathbb{P}(\theta_{0}\in CS)\geq 1-\alpha$, it follows that
@@ B.3 Proof of Corollary 1
Since for large enough sample size $n$, $||I_{n}(\tilde{\theta})-\tilde{\theta}||^{2}>||E_{n}(\tilde{\theta})-\tilde{\theta}||^{2}$, the chain of inequalities in 23 is well-defined. Therefore, we can find values for the tuning parameter $\sigma_{n}^{2}$ for all $n$’s that are sufficiently large such that a point $\tilde{\theta}$ in the complement of $CS$ will be labelled $-1$ by the SVM classifier.
Combining this with Theorem 1, it follows that when the sample size is large enough, by taking the values in the intersection of 22 and 23 for the tuning parameter $\sigma^{2}$, we have that a point $\theta\in\Theta$ belongs to $CS$ if and only if the SVM classifier under such tuning classifies it to have label $1$. ∎